THE VIBRATION-ROTATION EMISSION SPECTRA OF HgH_{2} AND HgD_{2}

\underline{ALIREZA SHAYESTEH}, SHANSHAN YU, PETER F. BERNATH, Department of Chemistry, University of Waterloo, Waterloo, ON, N2L 3G1, Canada.

Gaseous HgH_{2} and HgD_{2} were observed for the first time in a discharge of mercury vapor and molecular hydrogen at room temperature. The emission spectra were recorded in the 1200 − 2200 cm^{−1} spectral region using a Fourier transform spectrometer with 0.01 cm^{−1} resolution. The antisymmetric stretching fundamental bands of HgH_{2} and HgD_{2} were detected at 1912.8 cm^{−1} and 1375.8 cm^{−1}, respectively, and lines from the six naturally abundant isotopes of mercury were completely resolved. The 002 − 001 and 011 − 010 hot bands were also analyzed for both molecules. Unlike CdH_{2} and ZnH_{2}, no perturbation was observed in the HgH_{2} bands. The molecular constants and their isotopic ratios will be presented.